SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of Friday April 16 Machine Backgrounds meeting.

1. A. Valassi gave a status report on the miniTPC note describing Jan98 data.it will be ~15 pages long, half on description of apparatus, setup and offline software, half on results based on jan31 data : current scan and Neg experiment. It should be ready at the end of next week. A note for Jan-feb 1999 is awaiting for final MC release.

A BABAR note has been written for S. Petrak and A. Mc Kemey on the calorimeter module tests in Jan-feb 1999 and will be circulated shortly.

2. Report on action items from the BABAR/PEP-II weekly meeting

see the action item lists)

The SVT, EMC and DCH groups were asked to provide thresholds for beam startup. The proposed threshold were respectively 10 mrad/s , 0.1 rad/h, 0.5 rad./h. the EMC threshold of 0.1 rad/h corresponds to q diode current of 0.5 nA. Tom Mattison expresed doubts that this threshold be achievable given the temperature drifts.

Gerry Abrams reported on the work going on to define more states in the BABAR-PEP-II state machine (see attached mail). He also mentioned the need for an EPICS expert in our group to help

  1. Status of the SVTRAD board. Tim mentioned installation was progressing well
  2. Status of the analog system Tom mentioned that the boards have been sent out and should be ready next week. Thomas Schietinger gave the status of DCH and EMC diodes installation. They are in !

5. Review of background sensors . Terry gave an update of the table with the many EPICS addresses being

now listed. Uli stressed once more PEP-II requirements for these signals. The feeling was that all these requirements were satisfied for the set of sensors used in the protection devices (namely the PIN diodes) and may be a few extra signals, while it would take one week or two of beam expêrience to fully implement the rest .to the same level. Terry will mark this distinction in her list. Everybody providing a sensor should foprward to Terry documentation describing this sensor and calibration asap.

  1. Software week. The agenda of the Friday morning sessions devoted to background issues was discussed. The following proposal emerged and was sent to Nando Ferroni.

7. Report on background effects in EMC

Michael George showed the variation of resolution as function of background. It seems difficult to recover resolution in presence of background by tuning the clustering cuts.

Christos Touramanis discussed material posted on the hypernews. Degradation of width of pio peaks with PRV0 from 7 to 10 MeV.

Jean-Pierre Lees had material on the same topic , also posted on the hypernews. A ~60 MeV cut and a cut on pi0 energy above 300 MeV are needed to recover clean pi0 with PRV0 background.